GreenHab Report – December 21st

Greenhab Report
David Murray
21 Dec, 2017
Environmental control: (Choose which is appropriate and explain further if needed)
Working Hour: 14:06
Inside temp at working hour: 18 C
Outside temp during working hours: -0.5 C
Inside temperature high: 20 C
Inside temperature low: 14 C
Inside humidity: 16%
Functioning nominally,
Turns on at 16 C and off at 20 C
Not functional
Shade Cloth On/Off
The greenhab lights are on from 17:00 to 24:00 every night
Average temperature:
Changes to Crops
Morning Research Observations
Daily Water Usage for Crops
2 gallons
Time(s) of Watering Crops
We had snow today so the greenhab stayed relatively cool with a high of 20 C, the same temperature that the heater shuts off at. With a limited water supply, the cool temperatures were welcome as I was able to use 1/3 of the water that I normally do on the plants. No signs of the seeds sprouting but I wouldn’t expect it at this point anyway. Tomorrow I’ll transplant more tomatoes and select plants to use as the control and test subject of the Moringa fertilizer experiment. Because the duration of this experiment continues into the next crew, I’d like to set the time for reapplication around a week after the next crew arrives so they have time to get settled and comfortable.
Support/Supplies Needed

Operations Report – December 21st

Crew 185
Operations Report Thursday 12/21/2017
SOL: 3
Name of person filing report: A.Passeron
Non-nominal systems:
-Spacesuit N°3 :
The ventilation system stopped once again during the engineering EVA. I opened the backpack at 3:00pm and tried to find a faulty connection with the ohmmeter. All the wires are well connected. The faulty connection may thus be in the outlet itself. Still, the backpack can be charged by slightly moving the plug. The ventilation system is operational. Tomorrow, I will check the running time of the fully-charged spacesuit to test whether or not the battery is defective.
– Outer window of front airlock and central window of living area:
Windows covered with bubblewrap to insulate against cold.
– Water heater :
The water heater is plugged. The front screen displays 061 to 063. We haven’t touched anything except the plug yet.
At 8:20am, I went to the shower and turned the hot water on all the way for 1 minute but water remained cold.
In the living area, water does not flow into the sink when hot is turned on.
– Shower-head :
Shower-head leaks when turned on.
Generator (hours run):
Generator turned off at 8:00 am (SOC at 78%)
Generator turned on at 8:55am because of snowy weather (SOC at 69%)
Solar— SOC @ 5:05pm : 100%
Diesel – 80%
Propane – 56%
Ethanol Free Gasoline (5 Gallon containers for ATV) – 3.5 Gallons
Water (trailer) – 50 gallons
Water (static) – 320 gallons
Trailer to Static Pump used – No
Static to Loft Pump used – Yes
Water Meter: 50 gallons
Toilet tank emptied: Yes
ATVs Used: No
Oil Added? No
ATV Fuel Used:
# Hours the ATVs were used today: 00:00 hours
Notes on ATVs: ATVs were nominal.
Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Opportunity rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Curiosity rover used:  No
Beginning charge:
Ending charge:
Currently charging:
HabCar used and why, where?  No
Summary of internet: All Nominal
Summary of suits and radios: The outlet of spacesuit N°3 has a faulty connection. Checking the battery lifespan tomorrow.
Summary of Hab operations: Front airlock and living area windows are covered with bubblewrap. Shower-head leaks when turned on. Still trying to operate the water heater.
Summary of GreenHab operations: All nominal
Summary of ScienceDome operations: All nominal
Summary of RAMM operations: Not Operational
Summary of health and safety issues: Crew is Healthy
Questions, concerns and requests to Mission Support:

Journalist Report – December 21st

Journalist Report 21 Dec, 2017
[Sol 3]
Crew 185 had to readjust some of our plans today. After breakfast we noticed that it was snowing outside. At first, it was light and didnt stick to the ground. When Arno, Tibo, and John went to complete the engineering check the snow began to stick and visibility became very limited. Unfortuanetly we had to cancel our EVA because visibility was crucial for the task we had assigned. Tibo still got a chance to test the X-1 space suit in the snow but the expedition was very limited. After the three returned from their engineering check we drew a picture as a group to increase team cohesion.
For the rest of the night we need to make dinner, complete psychological tests, and finish our reports. The snow is no longer falling and only trace amounts remain on the ground. We’re placing an emphasis on stress reduction and leisure for the rest of the night. Tomorrow, we hope for more agreeable weather and an opportunity to complete our EVA.
Later nerds,
David Murray – Crew Biologist and Greenhab Officer
Copyright © The Mars Society. All rights reserved. | Main Site